# FIPS 204-Compatible Threshold ML-DSA via Masked Lagrange Reconstruction

### v4

@@ 4.7.3 Summary of UC Security
Table 4: Security profile comparison. P3+ = 1 signer step + 1 server round.
@@ 5.3 Performance Results
Table 5: Signing performance across threshold configurations (ML-DSA-65, Rust implementation, Intel i7-12700H). Values are averages over 5 independent runs, each with 5 trials per configuration. Due to rejection sampling variance, P2 exhibits high variance (up to $14\times$ between runs).
@@ 5.4 Protocol Profile Comparison
Profile P1 (TEE) is fastest at 5.8 ms but requires a trusted execution environment, making it suitable for HSM-based deployments. Profile P2 (5-round MPC) achieves 21.5 ms online time for 3-of-5 through parallel EdaBit-based A2B conversion and constant-depth AND tree for comparison; the 5-round structure merges commit-then-open (Rounds 3–4 $\to$ Round 2) and parallelizes comparison (Rounds 6–7 $\to$ Round 4). Profile P3+ provides 22.1 ms latency with FROST-like workflow: signers precompute nonces offline (2.5 KB per set) and respond within a time window (default: 5 min), requiring only a single user interaction. With GC precomputation, online 2PC evaluation drops to $\approx$12 ms.
@@ 5.5 Comparison with Prior Work
Table 7: Comparison with prior work. Times for prior schemes are from their original papers; our times are from our Rust implementation (ML-DSA-65, N=5N\!=\!5, T=3T\!=\!3). Our $\approx$21–45% success rate is comparable to ML-DSA’s expected $\approx$20–25%. P3+ = 1 signer step + 1 server round.
@@
The dPN25 compact scheme [dPN25] uses short-coefficient LSSS with binary reconstruction coefficients, achieving FIPS 204 compatibility but is limited to small thresholds ($T\leq 8$) due to the Ball-Çakan-Malkin bound [BCM21]. Noise flooding [BCDG21] masks large Lagrange coefficients with overwhelming statistical noise, enabling arbitrary thresholds but producing $5\times$ larger signatures that break FIPS 204 compatibility; its near-100% success rate comes from relaxed norm bounds. Ringtail [BMMSZ25] is based on a modified Dilithium variant (not standard ML-DSA) with custom parameters, achieving faster per-attempt time due to its 2-round structure.
Our scheme combines arbitrary thresholds with standard signature size and FIPS 204 verification compatibility. The observed $\approx$21–45% success rate is comparable to single-signer ML-DSA’s expected $\approx$20–25%, inherent to the FIPS 204 specification.
@@ 5.7 Scalability Analysis
Our benchmarks demonstrate practical scalability up to $T=32$ signers. Profile P1-TEE scales from 4 ms ($T=2$) to 60 ms ($T=32$), approximately linear in $T$. Profile P3+ scales from 17 ms ($T=2$) to 94 ms ($T=32$), maintaining sub-100ms latency throughout. Profile P2-MPC shows higher variance (21–194 ms) due to MPC overhead sensitivity to retry count, exceeding 100ms for $T\geq 16$.
For larger thresholds ($T>32$), the dominant cost becomes polynomial operations ($O(T\cdot n\cdot\log n)$ per signature). Extrapolating from our measurements, $T=64$ would require approximately 100–200 ms for P1-TEE and 150–250 ms for P3+.
@@ 5.8 Preprocessing Analysis
All profiles benefit from preprocessing. The Shamir nonce DKG (Section 3.4) adds one offline round of $O(|S|^{2})$ point-to-point messages ($\approx$3.8 KB each for ML-DSA-65) before each signing session. In practice, multiple nonce sharings can be precomputed during idle time and consumed on demand, amortizing the DKG cost across signing sessions.
@@
The main bottleneck remains MPC communication overhead (simulated as computation in our single-machine benchmark). In a distributed deployment, network latency would dominate, making preprocessing even more critical.
Deployment Latency Estimates.
Our benchmarks measure computation time on a single machine. In distributed deployments, end-to-end latency is approximately $t_{\mathsf{compute}}+\text{rounds}\times\text{RTT}$. On a LAN (1 ms RTT), P1 takes $\approx$10 ms and P3+ $\approx$19 ms for $(3,5)$-threshold. Over a WAN (20 ms RTT), P1 rises to $\approx$67 ms and P2 to $\approx$120 ms. Over intercontinental links (100 ms RTT), P2 reaches $\approx$520 ms, while P3+ remains under 220 ms due to its 2-round structure. Detailed estimates appear in the supplementary material. P3+’s reduced round count makes it the most network-friendly profile for geographically distributed deployments.
@@ 5.12 FIPS 204 Compliance Verification
Our “FIPS 204-Compatible” claim refers to signature format compatibility: threshold-generated signatures are byte-identical to single-signer ML-DSA signatures and pass any compliant FIPS 204 verifier without modification. Full NIST ACVP Known-Answer-Test (KAT) certification for production deployment would additionally require implementing the deterministic encoding functions specified in FIPS 204 Section 7.2.
@@ 6.1 Distributed Key Generation
Our protocol employs two distinct DKG protocols. The key DKG (executed once during setup) eliminates the trusted dealer assumption: each party contributes a random sharing of local secrets sampled from the short-coefficient distribution $\chi_{\eta}$, parties aggregate their shares, and Feldman commitments [Feldman87] provide robustness against malicious participants. The final secret $\mathbf{s}_{1}=\sum_{i}\mathbf{s}_{1}^{(i)}$ is unknown to any individual party. For lattice-specific VSS/DKG, see [FS24, GJKR07].
The nonce DKG (executed before each signing session) is our primary technical contribution. As described in Section 3.4, each party generates a degree-(T1)(T\!-\!1) polynomial whose higher-degree coefficients are sampled uniformly from $R_{q}^{\ell}$, producing a fresh Shamir sharing of the signing nonce. This structural choice enables statistical share privacy (Theorem 4.4). The nonce DKG can be preprocessed offline and amortized across multiple signing sessions; see the Supplementary Material for the full protocol details.
@@ 6.2 Optimistic Blame Protocol
  • Excessive aborts: More than $K=30$ consecutive failures (honest probability $(0.75)^{30}<10^{-4}$)
@@
This preserves privacy: all sensitive values stay inside the enclave; only cheater identities are output.
@@ 6.3 Profile P2: Fully Distributed Signing
For deployments where TEE trust is unacceptable, we provide a fully distributed protocol using secure multi-party computation. The core challenge is evaluating the r0-check predicate without reconstructing $c\mathbf{s}_{2}$ at any single point—since virtually all ML-DSA challenges $c$ are invertible in $R_{q}$, leaking $c\mathbf{s}_{2}$ would compromise the secret key.
@@ 5 Output
    Combiner-mediated commit-then-open (Rounds 3-4 $\to$ Round 2): All parties send $(Com_{i},masked_{i},rand_{i})$ to a designated combiner, who releases all openings simultaneously after collecting from all parties. SPDZ MAC check catches any tampering by the combiner.
@@
    Constant-depth comparison (Rounds 6-7 $\to$ Round 4): All 1536 coefficient comparisons are evaluated in parallel (Layer 1), followed by an AND-tree aggregation (Layer 2). Using Beaver triples, all $(d,e)$ values can be opened in a single round.
@@
The 5-round protocol is obtained from the 8-round base protocol (Appendix H) via two standard MPC composition steps. First, combiner-mediated commit-then-open merges sequential commit and open rounds (Rounds 3–4 $\to$ Round 2): parties send $(\mathsf{Com}_{i},\mathsf{masked}_{i},\mathsf{rand}_{i})$ to a designated combiner who releases all openings simultaneously; the combiner is an ordinary party, and SPDZ MAC checks detect any tampering, so this is UC-equivalent to a sequential broadcast [CLOS02]. Second, parallel Beaver-triple-based comparison merges the sequential comparison rounds (Rounds 6–7 $\to$ Round 4): all 1536 coefficient comparisons execute in parallel (Layer 1), followed by a single-round AND-tree aggregation (Layer 2). Each optimization replaces a sub-protocol with a functionally equivalent variant; by the UC composition theorem [Canetti01], the composed 5-round protocol UC-realizes the same ideal functionality. The hybrid argument in Appendix H ($\mathsf{H}_{0}$$\mathsf{H}_{6}$) carries over with the round assignment remapped, since each hybrid step depends on the functionality of each sub-protocol (SPDZ, edaBits, binary circuits), not on its round structure. The total security loss remains $\epsilon_{\mathsf{SPDZ}}+\epsilon_{\mathsf{edaBits}}<2^{-11}$.
@@
  • Bienstock et al. [BDEPT25]: 37–136 online rounds, honest majority
@@
  • Our P2: 5 online rounds, dishonest majority, arbitrary $T$
Profile P2 thus achieves dishonest-majority security with arbitrary thresholds in constant rounds.

### v5

@@ 4.7.3 Summary of UC Security
UC privacy level Statistical $\approx 2^{-11}$ Statistical
Table 4: Security profile comparison. P3+ = 1 signer step + 1 server round. UC privacy level for P2 refers to the r0-check simulation bound ($\epsilon_{\mathsf{SPDZ}}+\epsilon_{\mathsf{edaBits}}\leq 2^{-64}+2^{-12}<2^{-11}$); P1/P3+ achieve statistical privacy via the Shamir nonce DKG (no computational assumptions). P2 mask-hiding privacy requires $|S\setminus C|\geq 2$ (Lemma 4.15); P1 and P3+ achieve nonce share privacy with $|S\setminus C|=1$ ($|S|=T$).
@@ 5.2 Benchmark Methodology
Network Testing Setup.
For the $N=3$, $T=2$ configuration, we deployed parties across 3 physical nodes connected via OpenVPN (local area network). This setup validates protocol correctness and measures actual network communication overhead in a distributed environment. For larger configurations ($N\geq 5$), we benchmarked on a single machine; reported timings for these configurations represent computational cost only, with network latency projected analytically based on round count and representative RTT values (see the network latency table in the companion supplement document).
@@ 5.3 Performance Results
Table 5: Signing performance across threshold configurations (ML-DSA-65, Rust implementation, Intel i7-12700H). Values are averages over 5 independent runs, each with 5 trials per configuration. $N=3$: measured on 3-node OpenVPN testbed; $N\geq 5$: single-machine benchmarks (computational cost only; network overhead excluded). Due to rejection sampling variance, P2 exhibits high variance (up to $14\times$ between runs).
@@
Figure 2 extends the benchmark to $T=128$, showing that P3+ remains sub-1 s through $T=64$ and sub-3 s through $T=128$, while P1 and P2 exceed 10 s at T=96T=96128128 due to Lagrange polynomial evaluation and MPC simulation overhead, respectively.
[…]
@@ 5.4 Protocol Profile Comparison
Profile P1 (TEE) is fastest at 5.8 ms but requires a trusted execution environment, making it suitable for HSM-based deployments. Profile P2 (5-round MPC) achieves 21.5 ms online time for 3-of-5 through parallel EdaBit-based A2B conversion and constant-depth AND tree for comparison; the 5-round structure merges commit-then-open (Rounds 3–4 $\to$ Round 2) and parallelizes comparison (Rounds 6–7 $\to$ Round 4). Profile P3+ provides 22.1 ms latency with FROST-like workflow: signers precompute nonces offline (2.3 KB per set) and respond within a time window (default: 5 min), requiring only a single user interaction. With GC precomputation, online 2PC evaluation drops to $\approx$12 ms.
@@ 5.5 Comparison with Prior Work
Table 7: Comparison with prior work. Times for prior schemes are taken from their original papers on potentially different hardware configurations; direct cross-scheme comparison is indicative only. Our times are from our Rust implementation on an Intel i7-12700H (ML-DSA-65, N=5N\!=\!5, T=3T\!=\!3). Our $\approx$21–45% success rate is comparable to ML-DSA’s expected $\approx$20–25%. P3+ = 1 signer step + 1 server round.
@@
The dPN25 compact scheme [dPN25] uses short-coefficient LSSS with binary reconstruction coefficients, achieving FIPS 204 compatibility but is limited to small thresholds ($T\leq 8$) in practice (binomial reconstruction coefficients exceed $q$ at $T\geq 26$; asymptotically, the Ball-Çakan-Malkin bound [BCM21] gives $\Omega(T\log T)$ share size). Noise flooding [BCDG21] masks large Lagrange coefficients with overwhelming statistical noise, enabling arbitrary thresholds but producing $5\times$ larger signatures that break FIPS 204 compatibility; its near-100% success rate comes from relaxed norm bounds. Ringtail [BMMSZ25] is based on a modified Dilithium variant (not standard ML-DSA) with custom parameters, achieving faster per-attempt time due to its 2-round structure.
Our scheme combines arbitrary thresholds with standard signature size and FIPS 204 verification compatibility. The observed $\approx$21–45% success rate overlaps with single-signer ML-DSA’s expected $\approx$20–25%: the lower end (21%, $T=16$) is in line with single-signer, while the higher end (45%, $T=7$) reflects the Irwin-Hall distribution’s greater concentration near zero relative to a uniform nonce, placing more nonces within the acceptance region $\|\mathbf{z}\|_{\infty}<\gamma_{1}-\beta$ (see the Irwin-Hall success rate analysis in the companion supplement). The wide range (2× variation) is dominated by rejection-sampling variance, not systematic performance degradation.
@@ 5.7 Scalability Analysis
Our benchmarks demonstrate practical scalability up to $T=32$ signers. Profile P1-TEE scales from 4 ms ($T=2$) to 60 ms ($T=32$), approximately linear in $T$. Profile P3+ scales from 17 ms ($T=2$) to 94 ms ($T=32$), maintaining sub-100ms per-valid-signature latency throughout (Table 5; Figure 2 shows per-attempt times from a separate extended benchmark, which may be up to $2\times$ higher due to rejection-sampling retries). Profile P2-MPC shows higher variance (21–194 ms) due to MPC overhead sensitivity to retry count, exceeding 100ms for $T\geq 16$.
For larger thresholds ($T>32$), the dominant cost becomes polynomial operations ($O(T\cdot n\cdot\log n)$ per signature). Extended benchmarks (Figure 2) show P3+ achieves sub-1 s through $T=64$ and sub-3 s through $T=128$; P1 and P2 exceed 10 s at $T\geq 96$ due to Lagrange evaluation and MPC simulation overhead, respectively.
@@ 5.8 Preprocessing Analysis
All profiles benefit from preprocessing. The Shamir nonce DKG (Section 3.4) adds one offline round of $O(|S|^{2})$ point-to-point messages ($\approx$3.6 KB each for ML-DSA-65) before each signing session. In practice, multiple nonce sharings can be precomputed during idle time and consumed on demand, amortizing the DKG cost across signing sessions.
@@
The main bottleneck remains MPC communication overhead. For the $N=3$ OpenVPN testbed, measured round-trip latency averaged $\approx$2.1 ms (LAN setting). For larger configurations ($N\geq 5$) benchmarked on a single machine, network communication is not measured; in actual distributed deployments, network latency would dominate, making preprocessing even more critical.
Network Latency Analysis.
For single-machine benchmarks ($N\geq 5$), we project distributed deployment latency analytically as $t_{\mathsf{total}}\approx t_{\mathsf{compute}}+\text{rounds}\times\text{RTT}$, where $t_{\mathsf{compute}}$ is measured and RTT varies by deployment scenario. On a LAN (1 ms RTT), P1 takes $\approx$10 ms and P3+ $\approx$19 ms for $(3,5)$-threshold. Over a WAN (20 ms RTT), P1 rises to $\approx$67 ms and P2 to $\approx$120 ms. Over intercontinental links (100 ms RTT), P2 reaches $\approx$520 ms, while P3+ remains under 220 ms due to its 2-round structure. Detailed estimates appear in the supplementary material. P3+’s reduced round count makes it the most network-friendly profile for geographically distributed deployments.
@@ 5.12 FIPS 204 Compliance Verification
Our “FIPS 204-Compatible” claim refers to signature format compatibility: threshold-generated signatures are byte-identical to single-signer ML-DSA signatures and pass any compliant FIPS 204 verifier without modification. Full NIST ACVP Known-Answer-Test (KAT) certification for production deployment would additionally require the deterministic encoding functions specified in FIPS 204 Section 7.2.
@@ 6.1 Distributed Key Generation
Our protocol employs two distinct DKG protocols. The key DKG (executed once during setup) replaces a fully trusted dealer with a distributed protocol: each party contributes a random sharing of local secrets, parties aggregate their shares, and Feldman commitments [Feldman87] allow every participant to verify that all shares lie on a consistent polynomial without revealing the secret.
What Feldman VSS does and does not provide. Feldman VSS guarantees polynomial consistency: each party can verify that their share $\mathbf{s}_{1,h}=f(h)$ is consistent with the committed polynomial $\{g^{a_{0}},\ldots,g^{a_{T-1}}\}$. Crucially, this does not constrain the constant term $f(0)=\mathbf{s}_{1}$ to have short coefficients. The Shamir shares $f(h)$ for $h\geq 1$ lie in $\mathbb{Z}_{q}^{\ell}$ (large, as expected), which is fine for our protocol since the signing bound depends only on the reconstructed $\mathbf{s}_{1}$, not on individual share norms. The only constraint requiring smallness is that $\mathbf{s}_{1}=f(0)$ satisfies $\|\mathbf{s}_{1}\|_{\infty}\leq\eta=4$, which is needed for the rejection-sampling bound in the signature.
Why enforcing short $\mathbf{s}_{1}$ without trust is hard. To verifiably guarantee $\|\mathbf{s}_{1}\|_{\infty}\leq\eta$ in a trustless DKG, parties would need to produce a zero-knowledge proof that the committed polynomial has a small constant term. Current lattice-based ZK range proofs for $\ell_{\infty}$ balls in $R_{q}$ carry significant communication overhead; a practical treatment is an open problem (see Section 7). We note that this limitation is shared by all lattice-based threshold signature schemes in the literature—including Raccoon [RaccoTeam23] and Dilithium-based constructions—that require small-coefficient secrets.
Practical alternative. Profile P1 (Section 3.2) uses a TEE coordinator that runs ML-DSA key generation in a hardware-isolated environment and distributes Shamir shares directly. This provides the short-$\mathbf{s}_{1}$ guarantee at the cost of a hardware trust assumption (Intel SGX or equivalent), which is already required by P1’s security model. For deployments where hardware trust is acceptable—which covers a broad class of enterprise PKI scenarios—this is the recommended setup mechanism. For lattice-specific VSS/DKG, see [FS24, GJKR07].
The nonce DKG (executed before each signing session) is our primary technical contribution. As described in Section 3.4, each party generates a degree-(T1)(T\!-\!1) polynomial whose higher-degree coefficients (degree $\geq 1$) are sampled uniformly from $R_{q}^{\ell}$ and whose constant term is sampled from the bounded nonce range $\lfloor\gamma_{1}/|S|\rfloor,\lfloor\gamma_{1}/|S|\rfloor]^{n\ell}$, producing a fresh Shamir sharing of the signing nonce. This structural choice enables statistical share privacy (Theorem 4.7). The nonce DKG can be preprocessed offline and amortized across multiple signing sessions; see the Supplementary Material for the full protocol details.
@@ 6.2 Optimistic Blame Protocol
  • Excessive aborts: More than $K$ consecutive failures, where $K$ is set based on the deployment’s per-attempt success rate $p$: $K=\lceil\log(10^{-4})/\log(1-p)\rceil$. For $p\geq 25\%$ (all configurations with $T\leq 21$), $K=33$ suffices (honest probability $(0.75)^{33}<10^{-4}$); for worst-case $T=32$ ($p\approx 21\%$), set $K=48$ to maintain the $10^{-4}$ guarantee.
@@
Security Analysis.
The blame protocol achieves the following properties:
  • Privacy: All sensitive values ($\mathbf{f}_{h}(x)$, $\mathbf{s}_{1,i}$, mask seeds) are transmitted only to the TEE coordinator and never revealed to other parties. The TEE outputs only the identity of misbehaving parties, preserving statistical privacy of honest parties’ shares (Theorem 4.7 continues to hold even after blame execution).
  • Correctness: The TEE recomputes all protocol values deterministically from revealed inputs. A party is blamed if and only if its revealed inputs are inconsistent with its submitted protocol messages (e.g., $\mathbf{W}_{i}\neq\lambda_{i}\mathbf{A}\mathbf{y}_{i}+\mathbf{m}_{i}^{(w)}$), ensuring honest parties are never falsely accused.
  • Trigger threshold: The $K=33$ consecutive-failure threshold is chosen such that the probability of triggering blame due to honest rejection sampling alone is $(1-p)^{33}<10^{-4}$ for $p\geq 25\%$ (ML-DSA-65 with $T\leq 21$). For larger $T$, set $K=\lceil\log(10^{-4})/\log(1-p)\rceil$ to maintain the false-positive rate below $10^{-4}$.
  • UC composition: Under the TEE trust model ($\mathcal{F}_{\mathsf{TEE}}$ honest), the blame protocol UC-realizes an ideal functionality $\mathcal{F}_{\mathsf{Blame}}$ that takes inputs from all parties, identifies inconsistent parties via deterministic recomputation, and outputs cheater identities to all parties while keeping honest inputs private. Formal UC proof is deferred to the full version; the key observation is that the simulator can run the blame protocol honestly (since it has access to all corrupted parties’ inputs and can program the TEE’s output) and the view of honest parties consists only of the output (cheater identities), which is identically distributed in the real and ideal worlds.
Blame Protocol (Profile P2).
Profile P2 does not require a separate blame protocol: SPDZ MAC verification already provides identifiable abort. When a MAC check fails during the r0-check subprotocol, the SPDZ functionality $\mathcal{F}_{\mathsf{SPDZ}}$ outputs the identity of the party whose MAC is invalid (with probability $\geq 1-1/q\approx 1$ for $q=8380417$). The cheating party is immediately identified and excluded from future sessions. Note that the SPDZ MAC check catches input manipulation (wrong shares), but not wrong inputs submitted to $\mathcal{F}_{\mathsf{SPDZ}}$; a party that consistently submits out-of-range inputs to the edaBit protocol is caught by the cut-and-choose parameter $B=64$.
Blame Protocol (Profile P3+).
Profile P3+ uses 2PC, and blame attribution proceeds as follows when failures exceed the threshold $K$. Since CP1 does not hold key shares, blame is routed to CP2 (or a designated auditor):
  1. 1.
    CP2 requests each signer $i\in S$ to send: (a) their own nonce DKG polynomial $\mathbf{f}_{i}(x)$, and (b) all evaluations $\{\mathbf{f}_{j}(i)\}_{j\in S}$ they received during the DKG round.
  2. 2.
    CP2 reconstructs each party’s nonce share: $\mathbf{y}_{i}=\sum_{j\in S}\mathbf{f}_{j}(i)$.
  3. 3.
    Signer $i$ also sends to CP2 (over a confidential channel): their key share $\mathbf{s}_{1,i}$ and commitment randomness $r_{i}$. CP2 verifies: (a) $\mathsf{Com}_{i}=H(\texttt{"com"}\|\mathbf{y}_{i}\|\mathbf{w}_{i}\|r_{i})$ (commitment binding); (b) $\mathbf{w}_{i}=\mathbf{A}\mathbf{y}_{i}$ (nonce consistency); (c) $\mathbf{z}_{i}=\mathbf{y}_{i}+c\cdot\mathbf{s}_{1,i}$ (response correctness).
  4. 4.
    Inconsistent signers are identified and excluded.
Note on nonce privacy after blame. Revealing $\mathbf{f}_{i}(x)$ to CP2 discloses the constant term $\hat{\mathbf{y}}_{i}$ and all higher coefficients for the session under investigation, so per-session nonce privacy is intentionally forfeited during blame attribution. Theorem 4.7 (statistical nonce share privacy) holds for all sessions not under blame investigation; for the session under review, nonce privacy is deliberately sacrificed for accountability. Key privacy is maintained because $\mathbf{s}_{1,i}$ is sent only to CP2 (trusted auditor), not broadcast. After any blame invocation, a key-refresh step is recommended to limit exposure.
Remark 6.1 (Blame leakage across sessions).
Each blame invocation leaks at most $n\ell\cdot\log_{2}(2\lfloor\gamma_{1}/|S|\rfloor+1)\approx 25{,}600$ bits of information (the full session nonce polynomial $\mathbf{f}_{i}(x)$, with coefficients bounded in $\lfloor\gamma_{1}/|S|\rfloor,\lfloor\gamma_{1}/|S|\rfloor]$). Because each session uses freshly sampled higher polynomial coefficients, successive blame invocations on different sessions are statistically independent: the nonce polynomial for session $\tau$ reveals no information about session $\tau^{\prime}$. There is therefore no cumulative cross-session leakage from the nonce DKG component alone. However, once $\mathbf{s}_{1,i}$ is disclosed to CP2 (for blame verification), CP2 is effectively a key-holder. A mandatory key-refresh should follow the first blame invocation that requires revealing key material; subsequent blame invocations may reuse the refreshed key under the same analysis.
@@ 6.3 Profile P2: Fully Distributed Signing
For deployments where TEE trust is unacceptable, we provide a fully distributed protocol using secure multi-party computation. The core challenge is evaluating the r0-check predicate without reconstructing $c\mathbf{s}_{2}$ at any single point—since virtually all ML-DSA challenges $c$ are invertible in $R_{q}$ (a standard heuristic confirmed experimentally; see Remark Q.39), leaking $c\mathbf{s}_{2}$ would compromise the secret key.
@@ 5 Output
    Combiner-mediated commit-then-open (base-protocol Rounds 2–4 $\to$ 5-round Round 2): All parties send $(\mathsf{Com}_{i},\mathsf{masked}_{i},\mathsf{rand}_{i})$ to a designated combiner (one of the $N$ signing parties, potentially corrupted), who releases all openings simultaneously after collecting from all parties. SPDZ MAC checks catch any tampering by the combiner, since each share is authenticated with a global MAC key unknown to individual parties. This collapses three base-protocol rounds (masked-value exchange, commitment, and opening) into a single combiner-mediated round, saving two rounds.
@@
    Constant-depth comparison (base-protocol Rounds 6–7 $\to$ 5-round Round 4): All $nk=1536$ coefficient comparisons are evaluated in parallel using $1536$ Beaver triples; all $(d,e)$ openings are batched into a single network round (Round 4). After receiving opened $(d,e)$ values, each party locally computes comparison bits $b_{j}\in\{0,1\}$ and then locally aggregates $\langle\sum_{j}b_{j}\rangle$ via linear sharing (no additional network round). The AND of all 1536 comparison results reduces to: $\bigwedge_{j}b_{j}=1\iff\sum_{j}b_{j}=nk=1536$, which is a constant-depth (depth-1) check: opening $\langle\sum_{j}b_{j}\rangle$ in Round 5 (“Output”) and comparing to $1536$ constitutes a single network round with no sequential Beaver-triple dependencies.
@@
The 5-round protocol is obtained from the 8-round base protocol (Appendix H) via two standard MPC composition steps: Optimization 1 saves 2 rounds (collapsing three base-protocol rounds into one) and Optimization 2 saves 1 round (collapsing two parallel comparison rounds into one), giving 821=58-2-1=5 rounds total.
Optimization 1 (Combiner-mediated commit-then-open, base-protocol Rounds 2–4 $\to$ 5-round Round 2, saving 2 rounds): Parties send $(\mathsf{Com}_{i},\mathsf{masked}_{i},\mathsf{rand}_{i})$ to a designated combiner (one of the $N$ parties, treated as potentially corrupted) who releases all openings simultaneously. This is UC-equivalent to a sequential broadcast [CLOS02]: if the combiner withholds or alters any opening, the SPDZ MAC check — which authenticates each share with a global key unknown to the combiner — detects the tampering with probability $1-\epsilon_{\mathsf{SPDZ}}$. The simulator for this merged round constructs the same view as the 2-round version (send $\to$ open); leakage to the adversary is unchanged since only MAC-authenticated values are released. Confidentiality of the masked values $\mathsf{masked}_{i}$ from the combiner is protected by the pairwise PRF masks (Lemma 4.15), which ensure that even a corrupted combiner observing all $\mathsf{masked}_{i}$ cannot recover any party’s contribution, provided $|S\setminus C|\geq 2$.
Optimization 2 (Constant-depth comparison, base-protocol Rounds 6–7 $\to$ 5-round Round 4): All $nk=1536$ Beaver-triple comparisons are opened simultaneously in a single network round (Round 4). There are no sequential dependencies among the $1536$ Beaver triples: each triple $(d_{j},e_{j})=(\alpha_{j}\oplus\langle x_{j}\rangle,\,\beta_{j}\oplus\langle y_{j}\rangle)$ depends only on the party’s local share of $\langle x_{j}\rangle,\langle y_{j}\rangle$ and the preprocessed triple $(\langle\alpha_{j}\rangle,\langle\beta_{j}\rangle,\langle\gamma_{j}\rangle)$ with $\gamma_{j}=\alpha_{j}\wedge\beta_{j}$, none of which depend on other comparisons. After receiving all $(d_{j},e_{j})$ in Round 4, parties locally compute the comparison output bit $b_{j}=\gamma_{j}\oplus(d_{j}\wedge\beta_{j})\oplus(e_{j}\wedge\alpha_{j})\oplus(d_{j}\wedge e_{j})$ (standard Beaver-triple AND evaluation; the triple components $\alpha_{j},\beta_{j},\gamma_{j}$ are distinct from the output bit $b_{j}$) and aggregate $\langle B\rangle=\sum_{j}\langle b_{j}\rangle$ (purely local linear operation, no network round). The AND of all $1536$ comparison bits satisfies $\bigwedge_{j}b_{j}=1\iff B=nk$; opening $\langle B\rangle$ in Round 5 (“Output”) is a single network round. The AND aggregation therefore requires no sequential AND-gate rounds (the AND tree is replaced by a single linear aggregation + one opening round, with no Beaver triples).
By the UC composition theorem [Canetti01], each merged step preserves the ideal functionality’s input-output behavior and leakage profile. The simulator for the 5-round protocol is constructed by composing the simulators for each sub-protocol (SPDZ, edaBits, binary circuits) with the round assignment remapped; each simulator step depends on the functionality, not the round structure. The hybrid argument in Appendix H ($\mathsf{H}_{0}$$\mathsf{H}_{6}$) carries over directly. Since the edaBits functionality is used as a black box (its output distribution is indistinguishable from ideal, at cost $\epsilon_{\mathsf{edaBits}}$) before the SPDZ multiplication gate, there is no interaction between the two error terms, and the total security loss remains $\epsilon_{\mathsf{SPDZ}}+\epsilon_{\mathsf{edaBits}}<2^{-11}$.
@@
  • Bienstock et al. [BdCE25]: 23–79 online rounds (per [BdCE25]), honest majority
@@
  • Our P2: 5 online rounds, dishonest majority (unforgeability), arbitrary $T$
Profile P2 thus achieves dishonest-majority unforgeability with arbitrary thresholds in constant rounds. Note: commitment and r0-check mask privacy additionally requires $|S\setminus C|\geq 2$ (Lemma 4.15), meaning at most T2T-2 signing-set corruptions are tolerated for full mask privacy when $T\geq 3$.